Here are two vectors: a 2.00 m i 1.50 m j and ba 3.00 m i 4,00 m j. What are (a) the magnitude and (b) the angle (counterclockwise from the axis defined by I) of a What are (c) the magnitude and (d) the angle of b What are (e) the magnitude and (f) the angle of a b (g) the magnitude and (h) the angle of b a and (i) the magnitude and (j) the angle of a b (State your angle as a positive number (k) what is the angle between the directions of b a and a b (a) Number Units (b) Number Units (c) Number Units (d) Number Units (e) Number Units (f) Number Units (g) Number Units (h) Number Units (i) Number Units (j) Number Units (k) Number Units
Answer #1

a)

\underset{a}{\rightarrow} = 2 i - 1.5 j

magnitude = sqrt((2)2 + (- 1.5)2) = 2.5

b)

direction : \theta = 180 - tan-1(1.5/2) = 143.13 degrees

c)

\underset{b}{\rightarrow} = 3 i + 4 j

magnitude = sqrt((3)2 + (4)2) = 5

d)

direction : \theta = tan-1(4/3) = 53.12 degrees

e)

\underset{a}{\rightarrow}+ \underset{b}{\rightarrow} = 2 i - 1.5 j + 3 i + 4 j = 5 i + 2.5 j

magnitude : sqrt(52 + 2.52) = 5.6

f)

direction : \theta = tan-1(2.5/5) = 26.6 degrees